OpenStudy (anonymous):

Solve for b: 3(b+1)=4b-1 Explain how to do it please

OpenStudy (anonymous):

3(b+1)=4b-1 3b+3=4b-1 3b=4b-4 -b=-4 b=4

OpenStudy (anonymous):

expand 3(b+1) 3(b+1) = 3b + 3 3b + 3 = 4b - 1 b = 4